OpenStudy (badsheep):

A car accelerates from rest to a velocity of 5 meters/second in 4 seconds. What is its average acceleration over this period of time? Its acceleration is 1.25 meters/second. Its acceleration is 5 meters/second. Its acceleration is 0.8 meters/second. Its acceleration is 1.25 meters/second2. Its acceleration is 20 meters.

OpenStudy (theprofessor35):

5 m/s in 4 seconds m/s/s = m/s^2 5/4 = 1.25 1.25 m/s/s = Its acceleration is 1.25 meters/second2.

OpenStudy (theprofessor35):

Did that help you get the answer? :)

OpenStudy (badsheep):

thanks